SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

   
Prev Next


Date: January 11, 2008
Shift Taker: Chris West
-----------
General Features:
4-9 is slightly noisier because one diode is masked for each channel.

(theta,phi)
Warm crystals in hitmap          (52,63), (57,63), (56,64)
Warm crystals in 0.01 GeV hitmap (32,64), (17,43), (30,45), (32,46)
Warm crystals in 0.1 GeV hitmap  (32,64), (30,45), (32,46)
Warm crystals in 1 GeV hitmap      

Cool crystal  in hitmap          
Cool crystal  in 0.01 GeV hitmap 
Cool crystal  in 0.1 GeV hitmap  (29,85)
Cool crystal  in 1 GeV hitmap    (3,43)

78131-78132: Cosmics. OK.
78134-78138: OK.
78139: 1e-3 OOS, OOO in 0-5, 1e-3 OOS in 7-7.
78140: 2e-3 OOS in 1-6.
78141-78147: OK.
78153: 3e-2 OOS in 3-6.
78154: OK.


Crystal Spectra:
NTR.

Lightpulser:
27 dead channels, 0 channels out of time. Updated on: 05-January 2008 12:59, 19:33 PST



 

  Links

  • Shift Report Index
  • Back to the Shifter instructions